Find the square roots of 100 and 169 by the method of repeated subtraction.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

100
100–1=99
99–3=96
96–5=91
91–7=84
84–9=75
75–11=64
64–13=51
51–15=36
36–17=19
19–19=0
Here, we have performed subtraction ten times.∴√100=10

169
169–1=168
168–3=165
165–5=160
160–7=153
153–9=144
144–11=133
133–13=120
120–15=105
105–17=88
88–19=69
69–21=48
48–23=25
25–25=0
Here, we have performed subtraction thirteen times.∴√169=13

Thanks